Circular No. 6825

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 U1 (SOHO) C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports the discovery by D. A. Biesecker, University of Birmingham, of another comet in archival C3 coronagraphic data. The comet reached an estimated peak magnitude of 7.3 and never developed a tail. Measurements by Biesecker and St. Cyr have been reduced by G. V. Williams and are published in detail on MPEC 1998-D09, together with an orbit solution by B. G. Marsden.

1997 UT  α_{2000}  δ_{2000}
Oct. 28.134  13^{h}53.^{m}6  −16^{o}24′

V4334 SAGITTARII W. Liller, Instituto Isaac Newton, Viña del Mar, Chile; M. Janson,

Groningen University; H. W. Duerbeck, Münster University; and A. van Genderen, Leiden University, report that this object (cf. IAUC 6718) has faded by almost 2 mag since last November, indicating that the predicted dust formation (Duerbeck et al. 1997, A.J. 114, 1657) from the presumed final helium flash has begun. During 1997 Mar.–Nov., the star was between V = 10.75 and 11.0. Subsequent V magnitudes by Liller (L; 0.2-m telescope) and by Janson (J; ESO Dutch 0.9-m telescope): 1997 Nov. 21.02 UT, 10.76 (L); 1998 Feb. 9.37, 11.96 (L); 10.35, 11.90 (L); 14.38, 12.73 (J); 15.38, 12.75 (J); 17.36, 12.74 (J). The star has also become generally redder; between 1997 Sept. 14 and 1998 Feb. 15, U − B changed from +0.83 to +1.12, B − V from +1.66 to +1.51, V − R from +1.11 to +1.71, and V − i (Gunn i) from 2.08 to 2.86. Multicolor optical monitoring is being carried out, and infrared photometry is encouraged.

COMET 55P/TEMPEL-TUTTLE Total visual magnitude estimates: Feb. 1.00 UT, 8.3 (J. E. Bortle,

Stormville, NY, 20×80 binoculars); 3.3, 7.9 (S. J. O’Meara, Volcano, HI, 0.10-m refractor); 6.83, 8.3 (A. Pereira, Cabo da Roca, Portugal, 14×100 binoculars); 12.77, 8.4 (B. H. Granslo, Movatn, Oslo, Norway, 0.07-m refractor); 17.46, 8.4 (M. Mattiazzo, Wallaroo, S. Australia, 0.20-m reflector); 20.76, 8.7 (M. Meyer, Frauenstein, Germany, 0.25-m reflector).

1998 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green